# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ 4.4. The multilayered fluid-structure operator
[…]
$\enspace\mathbb{P}u-\mathbb{P}D_{\mathrm{fl}}\eta_{2}\in\mathrm{D}(A_{0})\enspace\text{and}\enspace(d_{1}-D_{\mathrm{s}}((P_{\mathrm{m}}\eta_{1})\mathrm{e}_{3}),d_{2}-D_{\mathrm{s}}((P_{\mathrm{m}}\eta_{2})\mathrm{e}_{3}))\in\mathrm{D}(A_{\mathrm{s}})\bigr{\}}.$
@@
(4.27) $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\begin{pmatrix}\mathbb{P}u\\ \eta_{1}\\ \eta_{2}\\ d_{1}\\ d_{2}\end{pmatrix}$ $=A_{\mathrm{mfs}}\begin{pmatrix}\mathbb{P}u\\ \eta_{1}\\ \eta_{2}\\ d_{1}\\ d_{2}\end{pmatrix}+\begin{pmatrix}\mathbb{P}f\\ 0\\ \overline{g}\\ 0\\ h\end{pmatrix},\enspace\text{for}\enspace t\in(0,T),\enspace\begin{pmatrix}\mathbb{P}u\\ \eta_{1}\\ \eta_{2}\\ d_{1}\\ d_{2}\end{pmatrix}(0)=\begin{pmatrix}\mathbb{P}u_{0}\\ \eta_{1,0}\\ \eta_{2,0}\\ d_{1,0}\\ d_{2,0}\end{pmatrix}.$
@@
[…]
Consider $q\in(1,\infty)$. Then there exists $\mu\geq 0$ such that the (shifted) multilayered fluid-structure operator $-A_{\mathrm{mfs}}+\mu$, with $A_{\mathrm{mfs}}$ as introduced in (4.25), is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}$ with $\mathcal{R}$-angle $\phi_{-A_{\mathrm{mfs}}+\mu}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. In particular, the operator $-A_{\mathrm{mfs}}+\mu$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}$.
@@
The second part of the assertion is an immediate consequence of the first one by the characterization of maximal $\mathrm{L}^{p}$-regularity in terms of $\mathcal{R}$-boundedness of the resolvents on UMD spaces such as the present $\mathrm{X}_{0}$ due to [59, Thm. 4.2], so we focus on the first part.
First, let us observe that the operator matrix $A_{\mathrm{mfs}}$ has a non-diagonal domain, which renders the analysis more complicated. In order to ease this, and to split the considerations for the coupling of the thin and the thick structural layer with the fluid, we will use a decoupling approach. More precisely, we will invoke a suitable similarity transform $S$, based on the lifting for the thick structural layer as explained in Subsection 4.3, and exploit that $\mathcal{R}$-sectoriality is preserved under similarity transforms, see for instance [16, Sec. 4.1].
@@
$S=\begin{pmatrix}\operatorname{Id}_{\mathrm{fs}}&0\\ T&\operatorname{Id}_{\mathrm{s}}\end{pmatrix}\enspace\text{and}\enspace S^{-1}=\begin{pmatrix}\operatorname{Id}_{\mathrm{fs}}&0\\ -T&\operatorname{Id}_{\mathrm{s}}\end{pmatrix},\enspace\text{with}\enspace T=\begin{pmatrix}0&-D_{\mathrm{s}}(P_{\mathrm{m}}(\cdot)\mathrm{e}_{3})&0\\ 0&0&-D_{\mathrm{s}}(P_{\mathrm{m}}(\cdot)\mathrm{e}_{3})\end{pmatrix}.$
@@
(4.29) $A_{\mathrm{fs}}$ $=\begin{pmatrix}A_{0}&0&-A_{0}\mathbb{P}D_{\mathrm{fl}}\\ 0&0&\operatorname{Id}\\ M_{s}^{-1}\gamma_{\mathrm{m}}N\Delta(\cdot)\cdot\nu&-M_{s}^{-1}P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}&M_{\mathrm{s}}^{-1}\Delta_{\mathrm{p}}\end{pmatrix},\enspace B=\begin{pmatrix}0&0\\ 0&0\\ M_{s}^{-1}\mathcal{K}&\delta M_{s}^{-1}\mathcal{K}\end{pmatrix}\enspace\text{and}\enspace$
@@
At this stage, let us recall from [41, Thm. 4.6] that the operator $A_{\mathrm{fs}}\colon\mathrm{D}(A_{\mathrm{fs}})\subset\mathrm{X}_{0}^{\mathrm{fs}}\to\mathrm{X}_{0}^{\mathrm{fs}}$ with domain
(4.30) $\mathrm{D}(A_{\mathrm{fs}})=\bigl{\{}(\mathbb{P}u,\eta_{1},\eta_{2})\in\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\mathrm{D}(A_{\mathrm{p}}):\mathbb{P}u-A_{0}\mathbb{P}D_{\mathrm{fl}}\eta_{2}\in\mathrm{D}(A_{0})\bigr{\}}$
has the property that $-A_{\mathrm{fs}}$ is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}^{\mathrm{fs}}=\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\times\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ with $\phi_{-A_{\mathrm{fs}}}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$.
@@
(4.31) $\tilde{A}_{\mathrm{mfs}}$ $=SA_{\mathrm{mfs}}S^{-1}=\begin{pmatrix}A_{\mathrm{fs}}-BT&B\\ TA_{\mathrm{fs}}-TBT+C-A_{\mathrm{s}}T&TB+A_{\mathrm{s}}\end{pmatrix},\enspace\text{with}\enspace$
@@
$BT$ $=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&-M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}&-\delta M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}\end{pmatrix},\enspace TB=\begin{pmatrix}0&0\\ -D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}&-\delta D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}\end{pmatrix}.$
@@
$TBT$ $=\begin{pmatrix}0&0&0\\ 0&D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}&\delta D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}\end{pmatrix}\enspace\text{and}\enspace A_{\mathrm{s}}T=\begin{pmatrix}0&0&-D_{\mathrm{s}}\\ 0&-L_{0}D_{\mathrm{s}}&-\delta L_{0}D_{\mathrm{s}}\end{pmatrix}.$
@@
(4.32) $\tilde{A}_{\mathrm{mfs}}=\begin{pmatrix}A_{\mathrm{fs}}-BT&B\\ \tilde{C}&TB+A_{\mathrm{s}}\end{pmatrix},\enspace\text{where}\enspace$
@@
$\tilde{C}$ $=\begin{pmatrix}0&0&0\\ -D_{\mathrm{s}}M_{s}^{-1}\gamma_{\mathrm{m}}N\Delta(\cdot)\cdot\nu&D_{\mathrm{s}}M_{s}^{-1}(P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}-\mathcal{K}D_{\mathrm{s}})&-D_{\mathrm{s}}M_{s}^{-1}(\Delta_{\mathrm{p}}+\delta\mathcal{K}D_{\mathrm{s}})\end{pmatrix}\enspace\text{and}\enspace$
@@
First, it readily follows from the result on $A_{\mathrm{fs}}$ that we recalled above together with Subsection 4.2(c) that the diagonal operator matrix $-\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}})$, with domain $\mathrm{D}(\tilde{A}_{\mathrm{mfs}})=\mathrm{D}(A_{\mathrm{fs}})\times\mathrm{D}(A_{\mathrm{s}})$, is $\mathcal{R}$-sectorial on $\mathrm{Y}_{0}$, with $\mathcal{R}$-angle strictly less than $\nicefrac{{\pi}}{{2}}$. Next, we consider $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}\colon\mathrm{D}(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})\subset\mathrm{Y}_{0}\to\mathrm{Y}_{0}$ defined by
(4.33) $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}=\begin{pmatrix}A_{\mathrm{fs}}&0\\ \tilde{C}&A_{\mathrm{s}}\end{pmatrix},\enspace\text{with}\enspace\mathrm{D}(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})=\mathrm{D}(\tilde{A}_{\mathrm{mfs}})=\mathrm{D}(A_{\mathrm{fs}})\times\mathrm{D}(A_{\mathrm{s}}).$
@@
(4.34) $\lambda-\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}=\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ -\tilde{C}&\lambda-A_{\mathrm{s}}\end{pmatrix}=\begin{pmatrix}\operatorname{Id}&0\\ -\tilde{C}R(\lambda,A_{\mathrm{fs}})&\operatorname{Id}\end{pmatrix}\cdot\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ 0&\lambda-A_{\mathrm{s}}\end{pmatrix}.$
@@
Thus, by the $\mathcal{R}$-sectoriality of $-\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}})$, we find that $-\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}$ is $\mathcal{R}$-sectorial on $\mathrm{Y}_{0}$ with $\phi_{-\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$.
@@
With regard to (4.32), it thus remains to discuss the part $\tilde{B}\coloneqq\begin{pmatrix}-BT&B\\ 0&TB\end{pmatrix}$, and to show that $\tilde{B}$ is a relatively bounded perturbation of $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}$, with arbitrarily small bound. In fact, as a preparation for the upcoming spectral theory, we will even verify that it is a relatively compact perturbation. For this purpose, let us first observe that the embedding $\mathrm{D}(L_{0})=\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\hookrightarrow\mathrm{W}^{2-\varepsilon,q}(\Omega)^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}$ is compact for all $\varepsilon>0$. In the following, we consider $\varepsilon>0$ sufficiently small so that $01-\nicefrac{{1}}{{q}}-\varepsilon>$ 0. By construction, it follows that $\mathcal{K}\colon\mathrm{W}^{2-\varepsilon,q}(\Omega_{\mathrm{s}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\to\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ is bounded. By Subsection 4.4(a), this yields the relative compactness of $B$. From Subsection 4.4(a), we also recall that $M_{s}$ is an automorphism on $\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$. On the other hand, we invoke the compactness of the embedding $\mathrm{W}^{4,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\hookrightarrow\mathrm{W}^{2-\nicefrac{{1}}{{q}},q}(\omega)$ joint with the boundedness of $D_{\mathrm{s}}\colon\mathrm{W}^{2-\nicefrac{{1}}{{q}},q}(\omega)\to\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$, $\mathcal{K}\colon\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}\to\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ and $M_{s}^{-1}\colon\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\to\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$. This implies the relative compactness of $BT$. With regard to the relative compactness of $TB$, we additionally invoke that (the extension of) $D_{\mathrm{s}}$ from (4.20) maps $\mathrm{L}^{q}(\omega)$ to $\mathrm{W}^{s,q}(\Omega_{\mathrm{s}})^{3}$ for all $s<\nicefrac{{1}}{{q}}$.
@@
In total, we have verified that $\tilde{B}$ is a relatively compact perturbation of $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}$. This already yields that $\tilde{B}$ is a relatively bounded perturbation of $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}$ with $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}$-bound zero, see [20, Lemma III.2.16]. Classical perturbation theory for $\mathcal{R}$-sectorial operators, see e. g. [34, Cor. 2], then yields that there exists $\mu\geq 0$ such that $-\tilde{A}_{\mathrm{mfs}}+\mu$ is $\mathcal{R}$-sectorial on $\mathrm{Y}_{0}$, with $\phi_{-\tilde{A}_{\mathrm{mfs}}+\mu}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. By the above argument on the preservation of the $\mathcal{R}$-sectoriality under similarity transforms, this completes the proof. ∎

### v2

@@ 5.4. The multilayered fluid-structure operator
(5.28) $\mathrm{D}(A_{\mathrm{mfs}})=\bigl\{$ $(\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})\in\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\mathrm{D}(A_{\mathrm{p}})\times\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}:$
$\kern 5.0pt\mathbb{P}u-\mathbb{P}D_{\mathrm{fl}}\eta_{2}\in\mathrm{D}(A_{0})\kern 5.0pt\text{and}\kern 5.0pt(d_{1}-D_{\mathrm{s}}((P_{\mathrm{m}}\eta_{1})\mathrm{e}_{3}),d_{2}-D_{\mathrm{s}}((P_{\mathrm{m}}\eta_{2})\mathrm{e}_{3}))\in\mathrm{D}(A_{\mathrm{s}})\bigr\}.$
@@
(5.29) $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\begin{pmatrix}\mathbb{P}u\\ \eta_{1}\\ \eta_{2}\\ d_{1}\\ d_{2}\end{pmatrix}$ $=A_{\mathrm{mfs}}\begin{pmatrix}\mathbb{P}u\\ \eta_{1}\\ \eta_{2}\\ d_{1}\\ d_{2}\end{pmatrix}+\begin{pmatrix}\mathbb{P}f\\ 0\\ \overline{g}\\ 0\\ h\end{pmatrix},\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0pt\begin{pmatrix}\mathbb{P}u\\ \eta_{1}\\ \eta_{2}\\ d_{1}\\ d_{2}\end{pmatrix}(0)=\begin{pmatrix}\mathbb{P}u_{0}\\ \eta_{1,0}\\ \eta_{2,0}\\ d_{1,0}\\ d_{2,0}\end{pmatrix}$.
@@
Proposition 5.14.
Consider $q\in(1,\infty)$. Then there exists $\mu\geq 0$ such that the (shifted) multilayered fluid-structure operator $A_{\mathrm{mfs}}+\mu$, with $A_{\mathrm{mfs}}$ as introduced in (5.27), is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}$ with $\mathcal{R}$-angle $\phi_{-A_{\mathrm{mfs}}+\mu}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. In particular, the operator $A_{\mathrm{mfs}}+\mu$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}$.
@@
The second part of the assertion is an immediate consequence of the first one by the characterization of maximal $\mathrm{L}^{p}$-regularity on UMD spaces, such as the present $\mathrm{X}_{0}$, as provided in Subsection 5.1, so we focus on the first part.
First, let us observe that the operator matrix $A_{\mathrm{mfs}}$ has a non-diagonal domain, which renders the analysis more complicated. In order to ease this, and to split the considerations for the coupling of the thin and the thick structural layer with the fluid, we will use a decoupling approach. More precisely, we will invoke a suitable similarity transform $S$, based on the lifting for the thick structural layer as explained in Subsection 5.3, and exploit that $\mathcal{R}$-sectoriality is preserved under similarity transforms, see Subsection 5.1(a).
@@
$S=\begin{pmatrix}\operatorname{Id}_{\mathrm{fs}}&0\\ T&\operatorname{Id}_{\mathrm{s}}\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0ptS^{-1}=\begin{pmatrix}\operatorname{Id}_{\mathrm{fs}}&0\\ -T&\operatorname{Id}_{\mathrm{s}}\end{pmatrix},\kern 5.0pt\text{with}\kern 5.0ptT=\begin{pmatrix}0&-D_{\mathrm{s}}(P_{\mathrm{m}}(\cdot)\mathrm{e}_{3})&0\\ 0&0&-D_{\mathrm{s}}(P_{\mathrm{m}}(\cdot)\mathrm{e}_{3})\end{pmatrix}$.
@@
(5.31) $A_{\mathrm{fs}}$ $=\begin{pmatrix}A_{0}&0&-A_{0}\mathbb{P}D_{\mathrm{fl}}\\ 0&0&\operatorname{Id}\\ M_{s}^{-1}\gamma_{\mathrm{m}}N\Delta(\cdot)\cdot\nu&-M_{s}^{-1}P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}&M_{\mathrm{s}}^{-1}\Delta_{\mathrm{p}}\end{pmatrix},\kern 5.0ptB=\begin{pmatrix}0&0\\ 0&0\\ M_{s}^{-1}\mathcal{K}&\delta M_{s}^{-1}\mathcal{K}\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0pt$
@@
At this stage, let us recall from [48, Thm. 4.6] that the operator $A_{\mathrm{fs}}\colon\mathrm{D}(A_{\mathrm{fs}})\subset\mathrm{X}_{0}^{\mathrm{fs}}\to\mathrm{X}_{0}^{\mathrm{fs}}$ with domain
(5.32) $\mathrm{D}(A_{\mathrm{fs}})=\bigl\{(\mathbb{P}u,\eta_{1},\eta_{2})\in\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\mathrm{D}(A_{\mathrm{p}}):\mathbb{P}u-\mathbb{P}D_{\mathrm{fl}}\eta_{2}\in\mathrm{D}(A_{0})\bigr\}$
has the property that $A_{\mathrm{fs}}$ is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}^{\mathrm{fs}}=\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\times\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ with $\phi_{-A_{\mathrm{fs}}}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$.
@@
(5.33) $\tilde{A}_{\mathrm{mfs}}$ $=SA_{\mathrm{mfs}}S^{-1}=\begin{pmatrix}A_{\mathrm{fs}}-BT&B\\ TA_{\mathrm{fs}}-TBT+C-A_{\mathrm{s}}T&TB+A_{\mathrm{s}}\end{pmatrix},\kern 5.0pt\text{with}\kern 5.0pt$
@@
$BT$ $=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&-M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}&-\delta M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}\end{pmatrix},\kern 5.0ptTB=\begin{pmatrix}0&0\\ -D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}&-\delta D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}\end{pmatrix}$.
@@
$TBT$ $=\begin{pmatrix}0&0&0\\ 0&D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}&\delta D_{\mathrm{s}}M_{s}^{-1}\mathcal{K}D_{\mathrm{s}}\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0ptA_{\mathrm{s}}T=\begin{pmatrix}0&0&-D_{\mathrm{s}}\\ 0&-L_{0}D_{\mathrm{s}}&-\delta L_{0}D_{\mathrm{s}}\end{pmatrix}$.
@@
(5.34) $\tilde{A}_{\mathrm{mfs}}=\begin{pmatrix}A_{\mathrm{fs}}-BT&B\\ \tilde{C}&TB+A_{\mathrm{s}}\end{pmatrix},\kern 5.0pt\text{where}\kern 5.0pt$
@@
$\tilde{C}$ $=\begin{pmatrix}0&0&0\\ -D_{\mathrm{s}}M_{s}^{-1}\gamma_{\mathrm{m}}N\Delta(\cdot)\cdot\nu&D_{\mathrm{s}}M_{s}^{-1}(P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}-\mathcal{K}D_{\mathrm{s}})&-D_{\mathrm{s}}M_{s}^{-1}(\Delta_{\mathrm{p}}+\delta\mathcal{K}D_{\mathrm{s}})\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0pt$
@@
First, it readily follows from the result on $A_{\mathrm{fs}}$ that we recalled above together with Subsection 5.2(c) that the diagonal operator matrix $\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}})$, with domain $\mathrm{D}(\tilde{A}_{\mathrm{mfs}})=\mathrm{D}(A_{\mathrm{fs}})\times\mathrm{D}(A_{\mathrm{s}})$, is $\mathcal{R}$-sectorial on $\mathrm{Y}_{0}$, with $\mathcal{R}$-angle strictly less than $\nicefrac{{\pi}}{{2}}$. Next, we consider the triangular block operator matrix $\tilde{A}_{\mathrm{mfs}}^{\triangle}\colon\mathrm{D}(\tilde{A}_{\mathrm{mfs}}^{\triangle})\subset\mathrm{Y}_{0}\to\mathrm{Y}_{0}$ defined by
(5.35) $\tilde{A}_{\mathrm{mfs}}^{\triangle}=\begin{pmatrix}A_{\mathrm{fs}}&0\\ \tilde{C}&A_{\mathrm{s}}\end{pmatrix},\kern 5.0pt\text{with}\kern 5.0pt\mathrm{D}(\tilde{A}_{\mathrm{mfs}}^{\triangle})=\mathrm{D}(\tilde{A}_{\mathrm{mfs}})=\mathrm{D}(A_{\mathrm{fs}})\times\mathrm{D}(A_{\mathrm{s}}).$
@@
(5.36) $\lambda-\tilde{A}_{\mathrm{mfs}}^{\triangle}=\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ -\tilde{C}&\lambda-A_{\mathrm{s}}\end{pmatrix}=\begin{pmatrix}\operatorname{Id}&0\\ -\tilde{C}R(\lambda,A_{\mathrm{fs}})&\operatorname{Id}\end{pmatrix}\cdot\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ 0&\lambda-A_{\mathrm{s}}\end{pmatrix}$.
@@
Thus, by the $\mathcal{R}$-sectoriality of $\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}})$, we find that $\tilde{A}_{\mathrm{mfs}}^{\triangle}$ is $\mathcal{R}$-sectorial on $\mathrm{Y}_{0}$ with $\phi_{-\tilde{A}_{\mathrm{mfs}}^{\triangle}}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$.
@@
With regard to (5.34), it thus remains to discuss the part $\tilde{B}\coloneqq\begin{pmatrix}-BT&B\\ 0&TB\end{pmatrix}$, and to show that $\tilde{B}$ is a relatively bounded perturbation of $\tilde{A}_{\mathrm{mfs}}^{\triangle}$, with arbitrarily small bound. In fact, as a preparation for the upcoming spectral theory, we will even verify that it is a relatively compact perturbation. For this purpose, let us first observe that the embedding $\mathrm{D}(L_{0})=\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\hookrightarrow\mathrm{W}^{2-\varepsilon,q}(\Omega)^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}$ is compact for all $\varepsilon>0$. In the following, we consider $\varepsilon>0$ sufficiently small so that $1-\nicefrac{{1}}{{q}}-\varepsilon>0$. By construction, it follows that $\mathcal{K}\colon\mathrm{W}^{2-\varepsilon,q}(\Omega_{\mathrm{s}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\to\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ is bounded. By Subsection 5.4(a), this yields the relative compactness of $B$. From Subsection 5.4(a), we also recall that $M_{s}$ is an automorphism on $\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$. On the other hand, we invoke the compactness of the embedding $\mathrm{W}^{4,q}(\omega)\cap\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\hookrightarrow\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$ joint with the boundedness of the operators
$D_{\mathrm{s}}$ $\colon\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3}\to\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3},\kern 5.0pt\mathcal{K}\colon\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}\to\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\kern 5.0pt\text{and}\kern 5.0pt$
$M_{s}^{-1}$ $\colon\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\to\mathrm{W}^{1-\nicefrac{{1}}{{q}}-\varepsilon,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega).$
This implies the relative compactness of $BT$. With regard to the relative compactness of $TB$, we additionally invoke that (the extension of) $D_{\mathrm{s}}$ from (5.22) maps $\mathrm{L}^{q}(\omega)$ to $\mathrm{W}^{s,q}(\Omega_{\mathrm{s}})^{3}$ for all $s<\nicefrac{{1}}{{q}}$.
@@
In total, we have verified that $\tilde{B}$ is a relatively compact perturbation of $\tilde{A}_{\mathrm{mfs}}^{\triangle}$. This already yields that $\tilde{B}$ is a relatively bounded perturbation of $\tilde{A}_{\mathrm{mfs}}^{\triangle}$ with $\tilde{A}_{\mathrm{mfs}}^{\triangle}$-bound zero, see [26, Lemma III.2.16]. Classical perturbation theory for $\mathcal{R}$-sectorial operators, see, e.g., [41, Cor. 2], then yields that there exists $\mu\geq 0$ such that $\tilde{A}_{\mathrm{mfs}}+\mu$ is $\mathcal{R}$-sectorial on $\mathrm{Y}_{0}$, with $\phi_{-\tilde{A}_{\mathrm{mfs}}+\mu}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. By the Subsection 5.1(a) on the preservation of the $\mathcal{R}$-sectoriality under similarity transforms, this completes the proof. ∎